# Towards Internet-Scale Training For Agents

### v1

@@ 4.1 Language Model Task Proposer
In the first stage, we scale the generation of web navigation tasks to 1M diverse websites using a Language Model Task Proposer. Shown in Figure 2, the task proposer serves two key functions in the pipeline: (1) filtering out harmful content and websites that cannot be safely annotated, and (2) proposing realistic web navigation tasks that a hypothetical user might want to accomplish.
Figure 2: Task proposal and filtering for 150k live websites. Starting from 1,000,000 websites, we employ a pretrained language model that marks sites as safe/unsafe for annotation, and assigns a realistic task that a hypothetical user might want to accomplish on each site. The task proposer aggressively filters out 85% of websites from the pipeline, resulting in 150k safe websites annotated with realistic tasks.
@@
We utilize pretrained and frozen language models that conform to a chat interface and accept a system prompt $\mathbf{x}_{\text{sys}}$, and a series of in-context examples through interleaved user and assistant prompts $\mathbf{x}_{\text{usr}}$ and $\mathbf{x}_{\text{ast}}$. The system prompt used for task generation is shown in Figure 2, and outlines all cases for which sites are considered unsafe for annotation. We consider the Llama 3.1 family of LLMs from Meta (Grattafiori et al., 2024; Touvron et al., 2023b; a), the GPT family of LLMs from OpenAI, and the Gemini family of LLMs from Google. Inference is served using vLLM (Kwon et al., 2023) for the Llama series of models. We employ a sampling temperature of 0.5, and a maximum budget of 64 newly generated tokens; all other parameters are kept as defaults in the OpenAI chat completions API, which is used to make inference calls to all LLMs.
@@
The goal of the task proposer is to accurately detect unsafe websites and generate realistic web navigation tasks when appropriate. We prompt the task proposer with the system prompt in Figure 2, a series of in-context examples (listed in Appendix E), and a final user prompt containing just the URL of the target website. We instruct the LLM via the system prompt to provide a task for the target website or to return “N/A” and mark the website as not suitable for annotation. This format produces a throughput of 20 websites per second for Llama 3.1 70B served on 16 GPUs with vLLM, processing 1M sites in 14 hours. The efficiency of stage one aids in scaling to large numbers of sites on the Internet, but we must not compromise safety and reliability for efficiency. To understand the trade-offs presented by our task proposal approach, we compare against typical human annotators at detecting safe websites for annotation and creating realistic agent tasks.
@@
Human Baseline 75% 0.71 0.84
Figure 3: Accuracy for detecting harmful sites. We select 100 website domains, where 50 are safe, and 50 are unsafe based on the criteria in Figure 2. Pretrained language models exceed the accuracy and recall of human annotators at detecting harmful sites that are unsuitable for training agents safely.
4.2 Improving Safety
Language models beat single pass non-expert annotators at detecting websites suitable for annotation. To evaluate detection performance, we employ the task proposer as a classifier and consider sites where the task proposer returns “N/A” as the positive class. We curate 50 safe and 50 unsafe domains, based on the filtering conditions outlined in the system prompt in Figure 2 (selected websites and their URLs are listed in Appendix E). We generate task proposals for each site and measure the accuracy, precision, and recall of our safety filter compared to human annotators. The annotators are asked to classify each site as suitable or unsuitable for annotation based on the website URL, and the criteria listed in the system prompt the same observations given to the task proposer to ensure a fair comparison. Results are presented in Table 3.
Understanding The Results.
Language models outperform human annotators by 29.3% in accuracy, 35.2% in precision, and 31.0% in recall at detecting harmful sites. While larger models like Gemini 1.5 Pro show the best overall accuracy, smaller models like Llama 3.1 70B display high recall with a minor decrease in accuracy. Recall matters most for safety filters, and these results suggest that Llama 3.1 70B is sufficient to detect most harmful sites with high confidence.
4.3 Improving Reliability
Language models are more reliable than human annotators at creating realistic web navigation tasks. To evaluate reliability, we measure the rate at which human workers are able to perform web navigation tasks generated by our pipeline. We curate a set of 100 safe website domains (different from the safety experiment, refer to Appendix E), generate task proposals using our pipeline, and measure the rate of self-reported task completion for human workers performing tasks. Workers start from the initial website URL in their browser and navigate pages using their mouse and keyboard while staying on the original site, reporting once the task is complete or once they believe the task is not feasible on the current site. We compare the feasibility rates for tasks generated by our pipeline and tasks written by human annotators given the criteria listed in Figure 2. Results are shown in Table 4.
Method Feasibility Rate
@@ 4.3 Improving Reliability
Human Baseline 54%
Figure 4: Expert feasibility of proposed tasks. We generate tasks for 100 safe websites (listed in Appendix E), and measure the completion rates of human workers attempting to perform the generated tasks in their browser. Language models exceed the performance of human annotators at creating feasible web navigation tasks for agents.
Understanding The Results.
Language models outperform human annotators by 64.8% at creating feasible web navigation tasks. Larger models like Gemini 1.5 Pro display the best feasibility rates, but the smaller model Llama 3.1 70B still outperforms the human annotators by 38.9%. To understand the relationship between the popularity of the site being annotated and the reliability of human-written tasks, we conduct an experiment in Figure 5 comparing PageRank values (Page et al., 1999) of sites according to the official June 2024 host-level web graph from The Common Crawl Foundation (2024), versus the feasibility rates of the proposed tasks from Table 4.
Although human annotators match the reliability of LLMs at creating feasible web navigation tasks for popular sites, LLMs outperform human annotators by 157.1% for less popular sites with low PageRank values. As obscurity increases, human annotators are less familiar with the sites, and the reliability of their task proposals decreases by 55.7%, whereas the reliability of tasks generated by LLMs remains relatively constant. This difference suggests that we should employ language models to ensure that task proposals are reliable as we begin to scale agents to vast numbers of sites on the Internet. However, where do we acquire this large and diverse set of websites to process for annotation?
Figure 5: Feasibility rates vs PageRank values. We visualize PageRank values, a proxy for the popularity of websites, versus the expert feasibility rates of proposed web tasks. Human-written tasks perform on par with LLMs for popular sites, but as target sites become less popular and annotators are less familiar with them, LLMs begin to outperform human annotators at creating feasible tasks for agents.
4.4 Scaling To 150,000 Sites
We propose to leverage open-source crawls of the internet for large-scale task generation. As of June, 2024, the web graph released by The Common Crawl Foundation (2024) contains more than 300 million unique hosts, which we adapt to a data source for agents. In particular, we sort hosts by their PageRank values, and select the top 1M sites for task generation. CommonCrawl is likely to contain many sites not suitable for annotation, and the experiments in Section 4.2 illustrate that the safety filter in the task proposer can detect and remove them effectively. In our configuration, task generation with Llama 3.1 70B takes 14 hours for 1M sites served with vLLM (Kwon et al., 2023) on two 8-GPU nodes. Sections 4.2 and 4.3 show Llama 3.1 70B outperforms human annotators in safety and reliability, and we can serve it locally at significantly reduced cost compared to proprietary LLMs with a marginal loss in quality. The distribution of tasks generated with Llama 3.1 70B for the top 1M sites in the CommonCrawl PageRank are visualized in Figure 6.
Understanding The Data.
The task proposer filters out 85% of the sites in CommonCrawl, resulting in 150k sites that can be safely assigned tasks for agents. Visualized in Figure 6, our distribution is denser than human-written tasks, has a wide coverage of real-world sites, and diverse categories of tasks. We automatically label task categories (procedure in Appendix E) and find that 89% of categories have fewer than the mean of 16.9 tasks per category. Top categories include news search, recipe search, product lookup, tutorial search, event schedules, health information, and many more. Refer to Appendix E for the top categories. Empowered by this large and diverse collection of tasks across the internet, we can start building internet-scale agents.
Figure 6: Distribution of 150k tasks. We compare the distribution of tasks generated by our pipeline (blue points) to the Mind2Web (Deng et al., 2023) dataset (orange points) via textual features extracted by a sentence embedding model, and projected in 2D with UMAP (McInnes et al., 2020). Our distribution is denser than human-written tasks, and has broad coverage of diverse real-world sites and tasks.
5 Internet-Scale Agents
In the next stage, we begin to scale LLM agents to diverse web navigation tasks across the web. Shown in Figure 7, we initialize a web browsing environment to the URL provided to the task proposer in Section 4, and run a language model agent to complete tasks by generating function calls in the Playwright API. For evaluation, current efforts typically use human-written constraints based on the final URL or page state (Zhou et al., 2024b; Koh et al., 2024a; Yao et al., 2023a; Drouin et al., 2024), but it can be difficult to scale these. Recall from Figure 5 that human annotators are less reliable for sites lower in the PageRank, where their familiarity is reduced. Results in section 4 showed that language models beat humans in safety and reliability for task generation. As we begin to scale agents to diverse websites from the internet, can we replace human-written criteria with language model judgments for efficient evaluation? Their robustness remains an important unresolved question, and current work only considers language model judges for a limited set of popular websites (He et al., 2024). We begin by validating the robustness of language models for evaluating diverse internet tasks.
@@ 5.1 Evaluation With Language Models
We model the process of evaluating trajectories from agents as a classification problem, where the goal is to estimate the probability $\mathbf{r}_{T}$ that a task $\mathbf{c}$ is solved by the final timestep. Conditioned on a system prompt $\mathbf{y}_{\text{sys}}$, a task $\mathbf{c}$, the history of previous actions $\mathbf{a}_{1},\cdots,\mathbf{a}_{T}$ and the final observation from a web browsing environment $\mathbf{s}_{T}$ formatted in text, we generate a completion using a language model and employ a function $f^{\;\text{text}\to\text{val}}(\cdot)$ to parse the estimated probability.
$\mathbf{r}_{T}=f^{\;\text{text}\to\text{val}}\big{(}\;\text{LLM}([\;\mathbf{y}_{\text{sys}},\mathbf{c},\mathbf{a}_{1},\cdots,\mathbf{a}_{T},\text{Enc}(\mathbf{s}_{T})\;])\;\big{)}$ (4)
Building on the sites used to measure reliability in Section 4.3, we conduct an experiment to measure the accuracy of language models to detect successful web navigation trajectories. We run agents on tasks generated by Llama 3.1 70B for the 100 sites in Section 4.3, and prompt language models to estimate the probability that tasks are solved $\mathbf{r}_{T}$ based on Equation 4. We then conduct human evaluations for the trajectories and manually assign binary success labels. Accuracy is calculated by applying a threshold to the predictions $\mathbf{r}_{T}>0.5$ to assign classes and tracking the rate at which the predictions agree with human labels. To understand robustness for sites of varying popularity, we report the accuracy of language models versus the PageRank of the corresponding sites. Similarly, to understand the ability of language models to judge their own uncertainty, we report their accuracy versus their prediction confidence, given by conf = $2\cdot|\mathbf{r}_{T}-1/2|$ (twice the total variation distance from the uniform distribution to the predicted distribution). Results are shown in Figure 8.
@@
Understanding The Results.
Language models are robust evaluators for web navigation tasks. Accuracy remains stable relative to PageRank values, suggesting that language models are effective for sites that typical human annotators are less familiar with. The best results are obtained with an evaluator based on GPT-4o, which attains an accuracy of $82.6\%$, compared to $81.7\%$ for Llama 3.1 70B, and $78.0\%$ for Gemini 1.5 Pro. Although the accuracy is robust to PageRank, the accuracy is highly informed by confidence. Language models show improved accuracy as their confidence improves, suggesting that they can effectively determine when their predictions are reliable. When considering predictions with conf = 1, the Llama 3.1 70B evaluator displays a compelling $93.1\%$ accuracy, 0.87 precision, and 0.82 recall for detecting successful trajectories. Now that we can efficiently and accurately judge trajectories, we can begin to scale language model agents to diverse internet tasks and track their success. Harnessing this judge, we can study the current abilities and shortcomings of language model agents spanning 150k websites.
Figure 9: Scaling LLM agents to 150k live sites. We run agents based on Llama 3.1 70B to complete tasks generated by our pipeline. We estimate success probabilities using a language model evaluator (left plot), and estimate probabilities agents are on the right track (right plot). 16.7% of tasks are estimated to be successful with conf = 1, and the spread of probabilities suggests data spans many difficulties.
@@ 5.2 Scaling To 150,000 Agents
We scale language model agents to 150k live sites in diverse domains across the Internet and attempt to complete 150k web navigation tasks generated by our pipeline. Shown in Figure 9, we evaluate the trajectories using a Llama 3.1 70B judge and run agents based on Llama 3.1 70B, selected because this model demonstrates high accuracy in Figure 8, and running currently available propriety models would be prohibitively expensive at this scale (see Appendix K for a cost analysis with different LLMs). We find that agents solve 16.7% of tasks with a model confidence of conf = 1. Furthermore, we observe that 35k tasks are judged to be on the right track with a confidence of conf = 1, suggesting that these could be solved if a larger computing budget was allocated. The spread along the x-axis in both plots in Figure 9 suggests that our tasks cover a broad range of difficulties, and working to solve them presents an opportunity to improve the capabilities of LLM agents. We observe that, when judging success, our evaluator tends to prefer binary predictions with high confidence values, suggesting that this subset of predictions is accurate based on Figure 8 results. Analyses for the agents that produced Figure 9 are presented in Appendix F.
Assembling the stages of the InSTA pipeline, we first generated tasks for 1M diverse sites across the Internet, we then dispatched language model agents to complete tasks for 150k sites marked as safe, and finally we evaluated the trajectories using a language model judge. Equipped with this large and diverse source of data, we seek to understand the utility of the data for training agents.
@@ 6 Training Agents
We compare the performance of agents trained on data from the InSTA pipeline to agents trained on human demonstrations from WebLINX (Lù et al., 2024) and Mind2Web (Deng et al., 2023), two recent and popular benchmarks for web navigation. Recent works that mix synthetic data with real data control the real data sampling probability in the batch $p_{\text{real}}$ independently from data size (Trabucco et al., 2024). We employ $p_{\text{real}}=0.5$ in few-shot experiments and $p_{\text{real}}=0.8$ otherwise. Shown in Figure 9, our data have a wide spread in performance, so we apply several filtering rules to select high-quality training data. First, we require the evaluator to return conf = 1 that the task was successfully completed, and that the agent was on the right track (this selects data where the actions are reliable, and directly caused the task to be solved). Second, we filter data where the trajectory contains at least three actions. Third, we remove data where the agent encountered any type of server error, was presented with a captcha, or was blocked at any point. These steps produce $7,463$ high-quality demonstrations in which agents successfully completed tasks on diverse websites. We sample 500 demonstrations uniformly at random from this pool to create a diverse test set, and employ the remaining $6,963$ demonstrations to train agents on a mix of real and synthetic data.
6.1 Improving Data-Efficiency
Figure 10: Data from InSTA improves efficiency. Language model agents trained on mixtures of our data and human demonstrations scale faster than agents trained on human data. In a setting with 32 human actions, adding our data improves Step Accuracy by +89.5% relative to human data for Mind2Web, and +122.1% relative to human data for WebLINX.
In a data-limited setting derived from WebLINX (Lù et al., 2024) and Mind2Web (Deng et al., 2023), agents trained on our data scale faster with increasing data size than human data alone. Without requiring laborious human annotations, the data produced by our pipeline leads to improvements on Mind2Web that range from +89.5% in Step Accuracy (the rate at which the correct element is selected and the correct action is performed on that element) with 32 human actions, to +77.5% with 64 human actions, +13.8% with 128 human actions, and +12.1% with 256 human actions. For WebLINX, our data improves by +122.1% with 32 human actions, and +24.6% with 64 human actions, and +6.2% for 128 human actions. Adding our data is comparable in performance gained to doubling the amount of human data from 32 to 64 actions. Performance on the original test sets for Mind2Web and WebLINX appears to saturate as the amount of human data increases, but these benchmark only test agent capabilities for a limited set of 150 popular sites.
6.2 Improving Generalization

### v2

@@ 4.1 Language Model Task Proposer
The key idea in stage one of the pipeline is a feedback loop, where a language model task proposer $\psi_{\text{LLM}}(\cdot)$ guides exploration on a website via an initial easy task. We then run a language model agent to explore the site, conditioned on the initial task, which produces an exploratory trajectory that is fed back to the task proposer. Conditioned on a trajectory that deeply explores the website, the task proposer then creates a harder, grounded task. This process is summarized as an equation.
$\mathbf{c}\sim\psi_{\text{LLM}}(\mathbf{c}|\mathbf{w},\tilde{\mathbf{c}},\underbrace{\mathbf{s}_{1},\mathbf{a}_{1},\cdots,\mathbf{s}_{T},\mathbf{a}_{T}}_{\text{trajectory}})\;\;\text{st}\;\;\mathbf{a}_{t}\sim\pi_{\text{LLM}}(\mathbf{a}_{t}|\mathbf{s}_{t},\tilde{\mathbf{c}})\;\;\mathbf{s}_{t+1}\sim\mathbb{P}(\mathbf{s}_{t+1}|\mathbf{s}_{t},\mathbf{a}_{t})$ (4)
In this equation, the website url is $\mathbf{w}$, the initial task is $\tilde{\mathbf{c}}$, the trajectory includes the states $\mathbf{s}_{t}$ and the actions $\mathbf{a}_{t}$, and the harder, grounded task is $\mathbf{c}$. Highlighted in Figure 2, we annotate 150k sites with tasks by scaling Equation 4, and release them on huggingface at: data-for-agents/insta-150k-v2.
Figure 2: Annotating 150k live sites with agentic tasks. Starting from 1,000,000 websites, we employ a pretrained language model that marks sites as safe/unsafe for annotation, and assigns a realistic task that a hypothetical user might want to accomplish on each site. The task proposer aggressively filters out 85% of websites from the pipeline, resulting in 150k safe websites annotated with realistic tasks.
@@
We utilize pretrained and frozen language models that conform to a chat interface and accept system, user, and assistant prompts. The task proposer system prompt is listed in Appendix 11, and details all cases for which sites are considered unsafe. We employ the Llama 3.1 family of LLMs from Meta (Grattafiori et al., 2024; Touvron et al., 2023b, a), the GPT family of LLMs from OpenAI, and the Gemini family of LLMs from Google. Inference is served using vLLM (Kwon et al., 2023) for the Llama series of models. We employ a sampling temperature of 0.5 and a maximum budget of 1024 new generated tokens, while all other parameters are kept as defaults in the OpenAI chat completions API, which is used to make inference calls to all LLMs.
@@
[…]
The design of the task proposer as a feedback process is important, and the full potential of this design will be realized in future work that trains agents with an on-policy reinforcement learning algorithm. In such future work, the task proposer can be used within the RL loop to generate incrementally harder tasks as agents learn. For this paper, we employ one loop of task generation.
Figure 3: Most frequent words in our tasks. This wordcloud shows the top 500 most frequent words in tasks from the training set of our official huggingface dataset. The size of each word corresponds to its frequency in the dataset. Our tasks span diverse categories and lexicon.
4.2 Safety & Reliability
Safety is critical when building autonomous agents. The internet contains significant amounts of content that should be removed from training data, in order to avoid agents learning harmful behaviors. To understand the robustness of our safety filter, we input 100 carefully selected websites to the task proposer, of which 50 contain harmful, or mature content. Table 2 reports the accuracy, precision, and recall of the safety filter on this data. For a variety of LLMs, the safety filter displays high accuracy—up to 97% of websites are correctly classified, and recall for detecting unsafe websites is as high as 1.0, suggesting that nearly all unsafe websites are detected by the safety filter.
@@ 4.2 Safety & Reliability
Table 1: The safety filter displays a high accuracy. We measure the accuracy, precision, and recall of the safety filter on a set of 100 websites, where 50 contain harmful, or mature content. Up to 97% of websites are correctly classified, and recall is as high as 1.0.
Method Verifiable Rate
@@
Table 2: Generated tasks are typically achievable. We measure the rate that human workers were able to complete and verify their completion of tasks produced by the task proposer for a set of 100 websites. Up to 89% of tasks are achievable, and verifiable.
Reliability is equally important for autonomous agents. Instructions should be followed faithfully by agents, and this requires training them with tasks that are achievable, and verifiable. Table 2 reports the rate that human workers were able to complete and verify their completion of tasks produced by the task proposer in its initial phase. Up to 89% of tasks are achievable, and verifiable according to the study, which suggests the pipeline is producing reliable tasks. Together with results in Section 6, it is likely that data from the InSTA pipeline leads to agents that follow instructions faithfully.
4.3 Scaling To 150,000 Websites
We leverage Common Crawl for task generation. As of May 2025, the latest web graph released by The Common Crawl Foundation (2025) contains more than 300M unique hosts, which we adapt into a data source for agents. In particular, we select the top 1M sites based on their PageRank values. Common Crawl likely contain many unsafe websites, and these are filtered out by the task proposer. Each phase of task generation requires 14 hours of compute time using two 8-GPU v100 machines, and filters to 150k safe websites annotated with tasks. Statistics of tasks are shown in Figure 3.
5 Internet-Scale Environment
By this point, we have reached our first goal—to improve coverage of real-world sites by annotating 150k diverse sites with challenging agentic tasks. To reach our second goal, and move beyond a static dataset, towards a dynamic internet-scale environment, we require a robust evaluator for these tasks. Evaluation presents a subtle challenge. The web evolves constantly, and daily changes in website content may invalidate a fixed ground truth reference solution. Driven by necessity, this environment must be evaluated by a model that judges whether an agent’s solution is correct, in the latest context.
@@ 5.1 Evaluation With Language Models
We model the process of evaluating trajectories from agents as a classification problem, where the goal is to estimate the probability $\mathbf{r}_{T}$ that a task $\mathbf{c}$ is solved, and generate $\mathbf{r}_{T}$ via next-token prediction, conditioned on a system prompt $\mathbf{x}_{\text{judge}}$, a task $\mathbf{c}$, and a trajectory $\mathbf{s}_{1},\mathbf{a}_{1},\cdots,\mathbf{s}_{T},\mathbf{a}_{T}$. The LLM is instructed to produce a reasoning trace that scores the agent on a scale from 0 to 1, and embed the score as JSON in a fenced code block. We employ a regex template that matches to the first code block in the response, and a JSON decoder to parse $\mathbf{r}_{T}$ from the response, given by $f^{\;\text{text}\to\text{val}}(\cdot)$.
$\mathbf{r}_{T}=f^{\;\text{text}\to\text{val}}(\;\text{LLM}(\;\mathbf{x}_{\text{judge}},\mathbf{c},\text{Enc}(\mathbf{s}_{1}),\mathbf{a}_{1},\cdots,\text{Enc}(\mathbf{s}_{T}),\mathbf{a}_{T}\;)\;)$ (5)
Verifying The Judge.
To understand the robustness of the judge, we measure its accuracy detecting successful trajectories that were annotated by human workers. We annotate 100 trajectories with binary success labels, and apply a threshold $\mathbf{r}_{T}>0.5$ to obtain binary predictions from the judge. Figure 5 reports the accuracy of the judge as a function of the PageRank values of sites, and as a function of the confidence of the judge, given by conf = $2\cdot|\mathbf{r}_{T}-1/2|$. For all LLMs tested, the judge shows a high accuracy, ranging from $78.0\%$ for Gemini 1.5 Pro, to $81.7\%$ for Llama 3.1 70B, and $82.6\%$ for GPT-4o. Accuracy is stable as PageRank falls, suggesting the judge is accurate for less popular sites that may not be well represented in the LLM’s training data. Shockingly, the confidence predicted by LLMs is highly interpretable, and correlates with accuracy to the point that trajectories where conf = 1 are classified with an accuracy up to $93.1\%$. The emergent robustness of the judge equips us to efficiently and accurately verify agent solutions on a dynamic internet.
@@
Figure 6: Statistics for our agent reasoning dataset. We conduct a large data collection experiment using our top checkpoint for Qwen3 1.7B. Our dataset has 2.2M screenshots, 2.2M reasoning traces for actions, and 150k traces for the judge. 50.0% of the trajectories are successful according to the judge, and have diverse ratings for efficiency and self-correction.
@@ 5.2 Scaling To 150,000 Agents
With our task proposer, agent, and judge driven by pretrained language models, we have all components needed to harness internet-scale data. We conduct a large data collection experiment, running language model agents to complete tasks on 150k websites from our official dataset, producing 2.2M screenshots and 2.2M reasoning traces for actions within 150k trajectories. The judge annotates these trajectories, producing 150k reasoning traces for evaluations, and leading to the statistics in Figure 6. For this experiment, we employ a fine-tuned Qwen3 1.7B as the agent (refer to the next section), and Qwen3 235B zero-shot as the judge. Data collection requires 1,200 v100 GPU hours, and costs \$521.55 based on current AWS spot instance pricing, a fraction of the budget that industry labs are spending towards agents. If you have the right data, this small budget is sufficient and no human annotations are required to build models that compete with frontier LLMs as agents.
@@ 6 Training Agents
Figure 7: InSTA unlocks great potential in small models. We train agents based on Qwen 3 1.7B using trajectories produced by a Qwen 3 235B data collection policy, and optionally filtered by a Qwen 3 235B judge (see Judge Filtered vs. Uniformly Sampled). We report success rates on a test set of 3,000 held-out websites and tasks. Before training, Qwen 3 1.7B has a zero-shot success rate of 11.5% according to a Qwen 3 235B judge, and we improve this by +45.3% absolute percentage points. Our top checkpoint outperforms the Qwen 3 235B data collection policy, and Llama 4 Maverick, a frontier LLM with 400B parameters, for which our model is 235 times smaller. Notably, filtering with a Qwen 3 235B judge leads to agents that improve according to independent secondary judges, including Gemini 2.5 Flash, Llama 4 Maverick, and Gpt 4.1 Nano, suggesting it generalizes well.
We’ve built an internet-scale data flywheel that produces trajectories annotated with scores from a judge that can help us train LLM agents. To understand the quality of data produced by the flywheel, we conduct a series of experiments training models on the data, and testing on popular benchmarks. These experiments focus on three main questions: (1) what is the impact of increasing data scale? (2) do agents transfer to new domains? (3) do agents scale with test-time compute?
6.1 Performance Improves With Data Scale
The recurring lesson in deep learning is that large-scale high-quality data wins, but researchers are struggling to materialize this promise for agents (Xue et al., 2025). Our paper aims to solve the data problem blocking researchers from materializing this promise, and this experiment provides a valuable signal that scaling high-quality data allows small models to compete with strong LLMs from top industry labs. To proceed, we collect 20k trajectories using a Qwen 3 235B data collection policy, annotated with scores from a Qwen 3 235B judge. We then train agents based on Qwen 3 1.7B with SFT on varying scales of the data. Results in Figure 7 show that performance improves with increasing data scale, and gains scale faster on data filtered by the judge. To filter the data, we select trajectories where Judge(Success) = 1. Our top checkpoint outperforms the Qwen 3 235B data collection policy, and beats Llama 4 Maverick, a frontier LLM with 400B parameters, for which our model is 235 times smaller. The trend in the figure suggests there is room to scale further.
Figure 8: Our agents zero-shot transfer to WebVoyager. With no additional training or specialized data, our checkpoints for Qwen 3 1.7B in Section 6.1 zero-shot transfer to the WebVoyager benchmark. Trends found on our test set appear to hold for WebVoyager as well, and our top checkpoints for Qwen 3 1.7B continue to match frontier LLMs in performance for three of four judges.
6.2 Agents Transfer To New Domains
Statistically correct evaluation for deep learning models requires a test dataset that does not overlap with the training dataset, but researchers do not agree on how to implement this guideline for agents. Recent works train agents on the same websites they test on (Murty et al., 2025; Zhou et al., 2024c; Putta et al., 2024), which may obfuscate progress (Xue et al., 2025). Our next experiment shows that we can implement a stronger train-test split. Agents trained with our data can zero-shot transfer to WebVoyager (He et al., 2024) without any data from the benchmark. Results in Figure 8 show our Qwen 3 1.7B checkpoint matching strong LLMs on WebVoyager (He et al., 2024) for three of four judges, confirming trends in Section 6.1. Our ability to zero-shot transfer relatively small agents to WebVoyager (He et al., 2024) makes it likely that our pipeline leads to capable agents.